# Generating crossmodal gene expression from cancer histopathology improves multimodal AI predictions

### v3

@@ Abstract
Emerging research has highlighted that artificial intelligence based multimodal fusion of digital pathology and transcriptomic features can improve cancer diagnosis (grading/subtyping) and prognosis (survival risk) prediction. However, such direct fusion for joint decision is impractical in real clinical settings, where histopathology is still the gold standard for diagnosis and transcriptomic tests are rarely requested, at least in the public healthcare system. With our novel diffusion based crossmodal generative AI model PathGen, we show that genomic expressions synthesized from digital histopathology jointly predicts cancer grading and patient survival risk with high accuracy (state-of-the-art performance), certainty (through conformal coverage guarantee) and interpretability (through distributed attention maps). PathGen code is available through GitHub at https://github.com/Samiran-Dey/PathGen for open use by the research community.
1 Introduction
@@
Figure 1: Data and methodology overview. (a-b) (i) Distribution of data across demographic categories (ii) Distribution of data across train - validation - calibration - test splits (iii) Number of genes in different gene groups. (c) Pipeline for the proposed methodology. Our novel diffusion-based crossmodal generative model, PathGen, synthesizes transcriptomic features from whole slide image patch embeddings obtained using a state-of-the-art foundation model from the whole slide images. Multimodal prediction is performed for both diagnosis (cancer grading) and prognosis (survival risk) using the synthesized transcriptomic data and histopathology images to predict the added value of transcriptomic data, for each patient. Uncertainty quantification provides a patient-level estimate of model reliability and provides clinicians with survival risk bounds and tumour grade prediction sets which are guaranteed to contain the ground truth with a specified probability. Distributed predictions also provide a window to understand the intra-tumour heterogeneity.
Classically histopathology has relied on microscopic assessment of H&\&&E stained tissue sections, often complemented by targeted panels of immunohistochemical and other special stains, to confirm and refine diagnoses. As technology has advanced, a broader range of molecular tests have become routine for assessing certain tumours. These include genetic analyses such as IDH mutation status, which is required to appropriately subtype gliomas according to the WHO Reuss [2023] and EGFR mutation status which is essential to guide therapy decisions in non-small cell lung cancers Wang and Fang [2014]. More recently, a range of transcriptomic tests have been developed, which show clear evidence for guiding treatment decisions in specific malignancies Sparano et al. [2018]. Following extensive clinical evaluation, several of these tests have been granted regulatory approval for clinical use Tsakiroglou et al. [2023]. However, guidelines, especially in public healthcare systems, strictly limit the application of these transcriptomic tests to a small minority of patients, largely due to their high financial cost and infrastructural requirements Harnan et al. [2019].
Recently, AI models employing a multimodal fusion of digital pathology and transcriptomic features, trained on large public datasets, like The Cancer Genomic Atlas (TCGA), have led to significant improvements in the predictive accuracy of both grading and survival risk estimation, compared to unimodal models employing digital H&\&&E slides alone Mobadersany et al. [2018a], Chen et al. [2020, 2021]. Though these powerful models have the potential to convey significant health benefits, they are developed in a setting disjoint from the current reality of the clinic. Translation of these models into most healthcare systems is infeasible at present, as the required transcriptomic data is very seldom collected. Future translation requires rigorous, costly and widespread clinical evaluation, to determine which patients would benefit from these models sufficiently to justify the cost of transcriptomic data.
@@
Here we introduce a novel, diffusion based crossmodal generative model, PathGen, that synthesizes transcriptomic data from digital images of H&\&&E stained slides. We develop our model on two publicly available multimodal data cohorts (glioma Network [2008, 2015] and renal cancer Network [2013]) from TCGA. We find that generated transcriptomic data are highly similar to the corresponding real transcriptomic data, in a hold-out validation set. We further demonstrate that the predictive performance of both diagnosis (cancer grading) and prognosis (survival risk) significantly improves when the synthesised transcriptomic data is combined with features learned from the corresponding digital pathology image, using a state-of-the-art ViT-based model through a co-attention mechanism.
@@
For practical utility, our model must be trusted by clinicians and attune to emerging regulation Commission [2021]. We thus provide transparency via detailed attention based heatmaps for both grading and survival risk prediction, allowing pathologists to visually affirm the clinical relevance of the regions contributing most towards the AI decision. We provide uncertainty quantification, via conformal analysis Angelopoulos and Bates [2021], to allow pathologist users, a measure of certainty in added value of the synthesized transcriptomic data for each patient. Finally, to evaluate algorithmic fairness, we apply our conformal uncertainty quantification to assess model performance analysis across various patient demographics. Our work builds on preliminary studies in crossmodal prediction Chen et al. [2024, 2020, 2021], to provide a road map for sustainable, safe and cost-effective translation of multimodal AI models supporting cancer diagnostics/prognostics, to the clinic.
Figure 2: Evaluation of synthesized transcriptomic data and explainability. (i) comparison between real and synthesized gene expression levels for different gene groups. (ii) plot of significant performance improvement on using synthesized transcriptomic data and not using them and closeness of using synthesized and real transcriptomics for gradation (AUC) and survival risk estimation (C Index). (iii) comparison of distributed prediction with non-distributed WSI prediction. (iv) comparison of co-attention maps generated using synthesized and real transcriptomic data. (v) percentage contribution of gene groups in co-attention.
2 Results
2.1 Dataset and study design
@@
Our experiments use publicly available data from TCGA  [^1] https://www.cancer.gov/tcga, specifically two independent cancer cohorts - Glioblastoma Multiforme and Brain Lower Grade Glioma (TCGA-GBMLGG) and Kidney Renal Clear Cell Carcinoma (TCGA-KIRC) Network [2013]. For TCGA-GBMLGG, 653 whole slide images (WSIs) coming from 532 cases are used to train, 69 WSIs from 56 cases to validate, 104 WSIs from 82 cases to calibrate and 86 WSIs from 75 cases to test (Figure 1.a.ii) of which 562 cases belong to TCGA-LGG Network [2015] and 183 cases belong to TCGA-GBM Network [2008]. For TCGA-KIRC Network [2013], we use 344 WSIs from 337 cases to train, 42 WSIs from 34 cases to validate, 48 WSIs from 45 cases to calibrate and 51 WSIs from 46 cases to test (Figure 1.b.ii). Cases with missing grade, survival data or transcriptomic data are excluded. The data distribution across various demographic groups like gender, age, censorship status, etc. varies for both cohorts (Figures 1.a.i, 1.b.i).
@@
Transcriptomic features are selected to match those employed in the successful multimodal models Chen et al. [2020, 2021]. As proposed previously Chen et al. [2021], transcriptomic data for each patient is divided into 6 broad categories of genes - tumour suppressor genes, oncogenes, protein kinases, cell differentiation markers, transcription factors, and cytokines and growth factors  [^2] https://github.com/mahmoodlab/MCAT/blob/master/datasets_csv_sig/signatures.csv. The number of available gene expression levels for each group are 84, 314, 498, 424, 1396 and 428 respectively for TCGA-GBMLGG resulting in a total of 3144 genes (Figure 1.a.iii). Whereas for TCGA-KIRC, the number of available gene expression levels for each group are 145, 522, 669, 681, 2366, and 990 respectively resulting in a total of 5373 genes (Figure 1.b.iii). When discussing gene expression levels throughout this paper, we refer to transcript abundance, rather than protein levels.
@@
Our methodology synthesizes transcriptomic data from WSIs using our diffusion based generative model, PathGen. Synthesized transcriptomes are then used for automated cancer gradation and survival risk estimation to assess whether transcriptomic tests should be requested (Figure 1.c). We use transcriptomic data (both real and synthesized) at two stages in the predictive model (MCAT_GR) - to co-attend the WSI patch embeddings and learn a combined feature map that is suitable to predict both grade and survival and to predict the survival risk of a patient in a learnable ensemble method along with the co-attended feature maps. We perform extensive experiments using synthesized transcriptomic data to decide on this model setting (supplementary Table 1). The model is trained end-to-end for performing both gradation and survival risk estimation together using the loss function, $L=\lambda\times L_{survival}+(1-\lambda)\times L_{grade}$, where $L_{survival}$ is the loss for survival risk estimation, $L_{grade}$ is the loss for gradation and $\lambda$ is the loss coefficient to balance the emphasis on survival risk estimation and gradation. Experiments show that at $\lambda=0.3$ the model performance is optimal with close and high scores for both gradation and survival risk estimation together for both cohorts (supplementary Figure 8 and supplementary Table 1). Thus for all results reported in the paper, we use $\lambda=0.3$. For uncertainty estimation using conformal prediction, we choose the error rate $\alpha=0.1$. Further details of model architecture, loss function and training methodology are provided in Methods.
Figure 3: Explainability analysis on TCGA-GBMLGG data. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomic data for chosen regions marked with different coloured rectangles.
2.2 Similarity evaluation of synthesized and real transcriptomic data
@@
The transcriptomic data synthesized by our model, PathGen, is compared to the real patient transcriptomes to assess similarity. The Spearman correlation between real and synthesized transcriptomic features comprising all genes evaluates to 0.71 for TCGA-GBMLGG ($p-value=1.18\times 10^{-67}$) and 0.72 ($p-value=3.65\times 10^{-69}$) for TCGA-KIRC, and the mean absolute error (MAE) evaluates to 0.1 for both cohorts (Figures 2.a.i, 2.b.i and supplementary Table 2). The gene expression levels for the gene groups show no significant differences between real and synthesized transcriptomes, for both cohorts (unpaired $t$-test $p$-value $>0.05$, supplementary Table 2). Though there exists a high correlation and low MAE between synthetic and real transcriptomic data, there is some variation across the gene sets (Figures 2.a.i, 2.b.i and supplementary Table 2). Notably, synthesized and real oncogene expression levels are highly correlated across both cohorts (Figures 2.a.i, 2.b.i). This is particularly beneficial for cancer grading or survival estimation as oncogenes when aberrantly active can serve as biomarkers that drive tumour progression, facilitating the prediction of cancer prognosis Shay [1999]. Thus we conclude that the synthesized transcriptomic data are highly correlated with the real transcriptomic data and we gain confidence towards using the former in the predictive pipeline.
2.3 Performance evaluation of synthesized transcriptomic data in multimodal prediction
2.3.1 Adding synthesized transcriptomic features with digital histopathology improves prediction
@@
We next analyse whether PathGen synthesised transcriptomes significantly add value to automated gradation and survival risk estimation, over and above whole slide images (WSIs) alone. Evaluation metrics used for grade prediction and survival risk estimation are the Area Under Curve (AUC) and the concordance index (C Index) respectively. Using WSIs alone we obtain an AUC of 0.823 (gradation) and C Index of 0.842 (survival risk) for TCGA-GBMLGG and an AUC of 0.714 (gradation) and a C Index of 0.671 (survival risk) for TCGA-KIRC (Figures 2.a.ii, 2.b.ii). The performance significantly improves when PathGen synthesized transcriptomics are used alongside WSIs (Wilcoxon rank-sum test $p$-value $<0.05$), resulting in an AUC of 0.890 (gradation) and C Index of 0.861 (survival risk) for TCGA-GBMLGG and an AUC of 0.773 (gradation) and C Index of 0.681 (survival risk) for TCGA-KIRC (Figures 2.a.ii, 2.b.ii). Thus, in the absence of real transcriptomic data, adding synthesized transcriptomic data significantly improves diagnostic and prognostic predictions compared to using only WSI features.
2.3.2 Comparison of predictive performance of synthesized and real transcriptomic data
@@
We next compare gradation and survival risk predictions based on WSIs and PathGen synthesized transcriptomic data to predictions based on WSIs and real transcriptomic data. Combining real transcriptomic data with WSIs yields an AUC of 0.907 (gradation) and C Index of 0.866 (survival risk) for TCGA-GBMLGG and an AUC of 0.778 (gradation) and C Index of 0.697 (survival) for TCGA-KIRC (Figures 2.a.ii, 2.b.ii) and the corresponding result using synthesized transcriptomic data is provided in Section 2.3.1. Remarkably, model performance using real transcriptomic data is not significantly different to using PathGen synthesized transcriptomes (Wilcoxon rank-sum test $p$-value $>$ 0.05) for both gradation and survival risk estimation, in both TCGA cohorts (supplementary Tables 4, 5). Thus, the transcriptomic data synthesized by our model, PathGen, performs comparably for gradation and survival risk estimation with no significant difference on average from real transcriptomic data.
2.3.3 Comparison with state-of-the-art methodologies
@@
We also compare the performance of our methodology with the state-of-the-art models for using real transcriptomic data and WSI patches. For TCGA-GBMLGG, the C Index evaluated for GSCNN Mobadersany et al. [2018b] is 0.781, DeepAttnMISL Yao et al. [2020] is 0.734, MCAT Chen et al. [2021] is 0.817, Pathomic fusion Chen et al. [2020] is 0.826 whereas our methodology yields a C Index of 0.866 using real transcriptomic data and a C Index of 0.861 using synthesized transcriptomic data for survival risk estimation. PathoGen-X Krishna et al. [2024] obtains a C-Index of 0.81 on TCGA-GBM cohort, however their inference stage is unimodal and only considers a section of patients (Grade IV). For gradation, Pathomic fusion Chen et al. [2020] has an AUC of 0.906 whereas our methodology has an AUC of 0.907 on real transcriptomic data and an AUC of 0.890 on synthesized transcriptomic data. For TCGA-KIRC, the C Index evaluated for Pathomic fusion Chen et al. [2020] is 0.720 whereas for our methodology is 0.697 on real transcriptomic data and 0.681 on synthesized transcriptomic data. However, Pathomic fusion uses ROIs annotated by expert pathologists for predicting grades, while we use transcriptomic features to co-attend to the WSI patch embeddings without requiring any expert intervention, thus reducing expert annotation overheads. We could only compare our model with Pathomic fusion for TCGA-GBMLGG gradation and TCGA-KIRC survival risk estimation as other models did not report the respective performances. Since our model performs better or at least comparably with the state-of-the-art models using real transcriptomic data, and we observe that the predictive performance between using real or synthesized transcriptomic data is very close with no significant difference for the test set population (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5), we may conclude that we achieve state-of-the-art performance using gene expression levels synthesized by our model, PathGen.
Figure 4: Explainability analysis on TCGA-KIRC data. (a) whole slide image (WSI) (b) intra-tumour gradation heterogeneity (c) intra-tumour survival heterogeneity (d) co-attention maps for real and synthesized transcriptomic data and WSI patches for different gene groups (e) comparative study of intra-tumour heterogeneity and corresponding co-attention maps obtained for prediction using synthesized transcriptomic data for chosen regions marked with different coloured rectangles.
2.4 Transparency evaluation for clinical interpretability
2.4.1 Visualization of intra-tumour heterogeneity using synthesized transcriptomic data
@@
To visualize intra-tumour heterogeneity, we use our trained model to predict grade and survival risk for individual WSI patches independently, in a distributed manner, allowing identification of regions of varied grade/survival risk importance within a single WSI (Figures 3.b, 4.b and supplementary Figures 10.b, 11.b) and survival estimates (Figures 3.c, 4.c and supplementary Figures 10.c, 11.c). Illustrative images are shown for patients of ground truth grade III, survival time 10.71 months, survival status alive (Figure 3); ground truth grade IV, survival time 26.08 months, survival status deceased (supplementary Figure 10) belonging to TCGA-GBMLGG, and for patients of ground truth grade I, survival time 33.99 months, survival status deceased (Figure 4); ground truth grade II, survival time 22.78 months, survival status alive (supplementary Figure 11) belonging to TCGA-KIRC. Some regions of interest are highlighted by bounding boxes of colours green, cyan and blue. It is evident that higher-grade regions are correlated to higher risks, as is expected (Figure 4.e). We find that considering WSIs at higher resolution helps to obtain more precise predictions for tissue regions, by obtaining distributed predictions at different magnification levels corresponding to a WSI, as is expected (supplementary Figure 9).
@@
Further, we consider the aggregate of the heterogenous (distributed) predictions and compare it with the grade and survival risk predicted by the model considering all WSI patches as input (non-distributed predictions). For TCGA-GBMLGG, the distributed performance is not significantly different to a one-time prediction with all WSI patches (ANOSIM test jointly over gradation and survival risk estimation $p$-value $>$ 0.05, Figure 2.a.iii). However, for TCGA-KIRC, performing distributed prediction improves the performance significantly (ANOSIM test jointly over gradation and survival risk estimation $p$-value $<$ 0.05, Figure 2.b.iii). Thus we see from the results that there is indeed a distributed variation of grade and risk when predicted locally for patches and mapped onto the WSIs as a heatmap, both within and across cohorts, which may provide a visual insight to the human expert regarding cancer heterogeneity.
2.4.2 Visualisation of how transcriptomic data map to histopathology images

### v4

@@ Abstract
Emerging research has highlighted that artificial intelligence-based multimodal fusion of digital pathology and transcriptomic features can improve cancer diagnosis (grading/subtyping) and prognosis (survival risk) prediction. However, such direct fusion is impractical in clinical settings, where histopathology remains the gold standard and transcriptomic tests are rarely requested in public healthcare. We experiment on two publicly available multimodal datasets, The Cancer Genomic Atlas and the Clinical Proteomic Tumor Analysis Consortium, spanning four independent cohorts: glioma-glioblastoma, renal, uterine, and breast, and observe significant performance gains in gradation and risk estimation ($p$-value <0.05<0.05) when incorporating synthesized transcriptomic data with WSIs. Also, predictions using synthesized features were statistically close to those obtained with real transcriptomic data ($p$-value >0.05>0.05), consistently across cohorts. Here we show that with our diffusion based crossmodal generative AI model, PathGen, gene expressions synthesized from digital histopathology jointly predict cancer grading and patient survival risk with high accuracy (state-of-the-art performance), certainty (through conformal coverage guarantee) and interpretability (through distributed co-attention maps). PathGen code is available on GitHub at https://github.com/Samiran-Dey/PathGen for open use.
INTRODUCTION
@@
Classically, histopathology has relied on microscopic assessment of H&E-stained tissue sections, often complemented by targeted panels of immunohistochemical and other special stains, to confirm and refine diagnoses. As technology has advanced, a broader range of molecular tests has become routine for assessing certain tumours. These include genetic analyses such as IDH mutation status, which is required to appropriately subtype gliomas according to the WHO Reuss [2023] and EGFR mutation status which is essential to guide therapy decisions in non-small cell lung cancers Wang and Fang [2014]. More recently, a range of transcriptomic tests have been developed, which show clear evidence for guiding treatment decisions in specific malignancies Sparano et al. [2018]. Following extensive clinical evaluation, several of these tests have been granted regulatory approval for clinical use Tsakiroglou et al. [2023]. However, guidelines, especially in public healthcare systems, strictly limit the application of these transcriptomic tests to a small minority of patients, largely due to their high financial cost and infrastructural requirements Harnan et al. [2019].
Recently, AI models employing a multimodal fusion of digital pathology and transcriptomic features, trained on large public datasets, like The Cancer Genomic Atlas (TCGA), have led to significant improvements in the predictive accuracy of both grading and survival risk estimation, compared to unimodal models employing digital H&E slides alone Mobadersany et al. [2018a], Chen et al. [2020, 2021]. Though these powerful models have the potential to convey substantial health benefits, they are developed in a setting disjoint from the current reality of the clinic. Translation of these models into most healthcare systems is infeasible at present, as the required transcriptomic data is very seldom collected. Future translation requires rigorous, costly and widespread clinical evaluation to determine which patients would benefit from these models sufficiently to justify the cost of transcriptomic data.
@@
In this work, we introduce a diffusion based crossmodal generative model, PathGen Dey et al. [2025], that synthesizes transcriptomic data from digital images of H&\&E-stained slides. We evaluate our model on two publicly available multimodal datasets: The Cancer Genomic Atlas (TCGA) Network [2013a] and the Clinical Proteomic Tumor Analysis Consortium (CPTAC) Edwards et al. [2015], using data from four independent cancer cohorts: glioma-glioblastoma, renal, uterine and breast. We find that the generated transcriptomic data are highly similar to the corresponding real transcriptomic data in a hold-out test set. We further demonstrate that the predictive performance of both diagnosis (cancer grading) and prognosis (survival risk) significantly improves when the synthesised transcriptomic data is combined with features learned from the corresponding digital pathology image, using a state-of-the-art ViT-based model through a co-attention mechanism. For practical utility, our model must be trusted by clinicians and attune to emerging regulation Commission [2021]. We thus provide transparency via detailed attention based heatmaps for both grading and survival risk prediction, allowing pathologists to visually affirm the clinical relevance of the regions contributing most towards the AI decision. We provide uncertainty quantification, via conformal analysis Angelopoulos and Bates [2021], to allow pathologist users a measure of certainty in the added value of the synthesized transcriptomic data for each patient. Finally, to evaluate algorithmic fairness, we apply our conformal uncertainty quantification to assess model performance analysis across various patient demographics. Our work builds on preliminary studies in crossmodal prediction Chen et al. [2024, 2020, 2021], to provide a road map for sustainable, safe and cost-effective translation of multimodal AI models supporting cancer diagnostics/prognostics, to the clinic.
RESULTS
Dataset and study design
@@ RESULTS
Our experiments use publicly available data from TCGA and CPTAC, specifically from four independent cancer cohorts - Glioblastoma Multiforme and Brain Lower Grade Glioma (TCGA-GBM, TCGA-LGG, CPTAC-GBM), Kidney Renal Clear Cell Carcinoma (TCGA-KIRC), Uterine Corpus Endometrial Carcinoma (TCGA-UCEC, CPTAC-UCEC), and Breast Invasive Carcinoma (TCGA-BRCA). For TCGA-GBMLGG, 912 whole slide images (WSIs) coming from 745 cases are used, of which 183 cases belong to TCGA-GBM Network [2008] and 562 cases belong to TCGA-LGG Network [2015]. We consider 485 WSIs from 462 cases for TCGA-KIRC Network [2013b], 294 WSIs from 267 cases for TCGA-UCEC Erickson et al. [2016], 1010 WSIs from 946 cases for TCGA-BRCA Lingle et al. [2016], 242 WSIs from 62 cases for CPTAC-GBM Consortium et al. , and 364 WSIs from 71 cases for CPTAC-UCEC National Cancer Institute Clinical Proteomic Tumor Analysis Consortium (2019) [CPTAC] for our experiments. The slides have been split into training, validation, calibration, and test sets (Supplementary Table LABEL:s-data_split). The models trained using the TCGA data have been evaluated on the corresponding CPTAC cohorts, without any further training or finetuning. Cases with missing grade, survival data or transcriptomic data were excluded. The data distribution across various demographic groups like gender, age, censorship status, etc. varies for all cohorts (Supplementary Table LABEL:s-data_demo).
@@
Transcriptomic features are selected to match those employed in the successful multimodal models Chen et al. [2020, 2021]. As proposed previously Chen et al. [2021], transcriptomic data for each patient is divided into 6 broad categories of genes - tumour suppressor genes, oncogenes, protein kinases, cell differentiation markers, transcription factors, and cytokines and growth factors. However, a gene may occur in multiple gene groups. The number of available gene expression levels for each gene group differs across cohorts but remains same for the TCGA and CPTAC datasets for the same cancer cohort (Supplementary Table LABEL:s-data_trans). When discussing gene expression levels throughout this paper, we refer to transcript abundance, rather than protein levels.
@@
Our methodology synthesizes transcriptomic data from WSIs using our diffusion based generative model, PathGen. Synthesized transcriptomes are then used for automated cancer gradation and survival risk estimation to assess whether transcriptomic tests should be performed (Figure 1.c). Transcriptomic data (both real and synthesized) is used at two stages in the predictive model (MCAT_GR) - to co-attend to the WSI patch embeddings and learn a combined feature map that is suitable to predict both grade and survival, and to predict the survival risk of a patient in a learnable ensemble method along with the co-attended feature maps. We perform extensive experiments using synthesized transcriptomic data to decide on this model setting (Supplementary Table LABEL:s-result_gen_use). The model is trained end-to-end for performing both gradation and survival risk estimation together using the loss function $L=\lambda\times L_{survival}+(1-\lambda)\times L_{grade}$, where $L_{survival}$ is the loss for survival risk estimation, $L_{grade}$ is the loss for gradation and $\lambda$ is the loss coefficient to balance the emphasis between survival risk estimation and gradation. Experiments show that at $\lambda=0.3$ the model performance is optimal with close and high scores for both gradation and survival risk estimation together across glioblastoma-glioma and renal cohorts (Supplementary Figure LABEL:s-fig_lambda and Supplementary Table LABEL:s-result_gen_use). Thus, for all results reported in the paper, we use $\lambda=0.3$. For uncertainty estimation using conformal prediction, we choose the error rate $\alpha=0.1$. Further details of model architecture, loss function and training methodology are provided in Methods.
Similarity evaluation of synthesized and real transcriptomic data
@@
The transcriptomic data synthesized by our model, PathGen, is compared to the real patient transcriptomes to assess similarity. The Spearman correlation coefficient for all gene groups between real and synthesized transcriptomic features evaluates to 0.713 for TCGA-GBMLGG, 0.717 for TCGA-KIRC, 0.436 for TCGA-UCEC, 0.642 for TCGA-BRCA, 0.662 for CPTAC-GBM, and 0.669 for CPTAC-UCEC (Figure 2.a.i and Supplementary Table LABEL:s-result_trans). Thus, showing highly significant correlation between real and synthesized transcriptomic data (Spearman correlation $p$-value <0.05<0.05). The corresponding normalized mean absolute error (nMAE) over all gene groups evaluates to 0.141, 0.160, 0.173, 0.155, 0.167 and 0.178 for TCGA-GBMLGG, TCGA-KIRC, TCGA-UCEC, TCGA-BRCA, CPTAC-GBM and CPTAC-UCEC respectively (Figure 2.a.i and Supplementary Table LABEL:s-result_trans). Though there exists a high correlation and low MAE between synthetic and real transcriptomic data, there are some variations in correlation across the gene sets and data cohorts (Figures 2.a.ii, 2.a.iii and Supplementary Table LABEL:s-result_trans). Notably, synthesized and real oncogene expression levels are highly correlated for glioblastoma-glioma, renal and breast cohorts (Figure 2.a.iii and Supplementary Table LABEL:s-result_trans). This is particularly beneficial for cancer grading or survival estimation because oncogenes when aberrantly active can serve as biomarkers that drive tumour progression, facilitating the prediction of cancer prognosis Shay [1999]. Thus we conclude that the synthesized transcriptomic data are highly correlated with the real transcriptomic data and we gain confidence towards using the former in the predictive pipeline.
Performance evaluation of synthesized transcriptomic data in multimodal prediction
@@
We next analyse whether PathGen synthesised transcriptomes significantly add value to automated gradation and survival risk estimation, over and above whole slide images (WSIs) alone. Evaluation metrics used for grade prediction and survival risk estimation are the Area Under Curve (AUC) and the concordance index (C Index) respectively. For survival risk estimation the C Index obtained using WSIs alone are 0.842, 0.671, 0.663, 0.603, 0.547, 0.518 for TCGA-GBMLGG, TCGA-KIRC, TCGA-UCEC, TCGA-BRCA, CPTAC-GBM and CPTAC-UCEC repectively (Figure 2.b.i and Supplementary Table LABEL:s-result_main). And for gradation, we obtain an AUC of 0.823 for TCGA-GBMLGG, 0.714 for TCGA-KIRC, 0.796 for TCGA-UCEC, 0.736 for CPTAC-GBM, and 0.508 for CPTAC-UCEC (Figure 2.b.ii and Supplementary Table LABEL:s-result_main). Since TCGA-BRCA does not report cancer grades in the clinical data, only survival risk estimation could be performed. The performance significantly improves when PathGen synthesized transcriptomics are used alongside WSIs (Wilcoxon rank-sum test $p$-value <0.05<0.05), resulting in an AUC of 0.890 and C Index of 0.861 for TCGA-GBMLGG, AUC of 0.773 and C Index of 0.681 for TCGA-KIRC, AUC of 0.821 and C Index of 0.673 for TCGA-UCEC, C Index of 0.720 for TCGA-BRCA, AUC of 0.865 and C Index of 0.565 for CPTAC-GBM, and AUC of 0.593 and C Index of 0.530 for CPTAC-UCEC (Figures 2.b.i, 2.b.ii and Supplementary Table LABEL:s-result_main). A significant improvement was also noted on evaluating jointly over survival and gradation for using synthesized transcriptomic data with WSIs for prediction (ANOSIM test $p$-value =0.001=0.001, Supplementary Table LABEL:s-result_main), consistent over datasets and cohorts. Thus, in the absence of real transcriptomic data, adding synthesized transcriptomic data significantly improves diagnostic and prognostic predictions compared to using only WSI features.
@@
We further compare gradation and survival risk predictions using WSIs and PathGen synthesized transcriptomic data to predictions using WSIs and real transcriptomic data. Combining real transcriptomic data with WSIs yields an AUC of 0.907 and C Index of 0.866 for TCGA-GBMLGG jointly, AUC of 0.778 and C Index of 0.697 for TCGA-KIRC, AUC of 0.828 and C Index of 0.680 for TCGA-UCEC, C Index of 0.720 for TCGA-BRCA, AUC of 0.863 and C Index of 0.564 for CPTAC-GBM, and AUC of 0.593 and C Index of 0.533 for CPTAC-UCEC and the corresponding result using synthesized transcriptomic data is provided above (Figures 2.b.i, 2.b.ii and Supplementary Table LABEL:s-result_main). Remarkably, model performance on using real transcriptomic data is not significantly different to using PathGen synthesized transcriptomes (Wilcoxon rank-sum test $p$-value >0.05>0.05 and ANOSIM test $p$-value >0.05>0.05) for gradation and survival risk estimation, both separately and jointly, in all considered TCGA and CPTAC cohorts (Supplementary Table LABEL:s-result_main). Thus, the transcriptomic data synthesized by our model, PathGen, consistently performs comparably for gradation and survival risk estimation with no significant difference on average from real transcriptomic data.
@@
We also compare the performance of our methodology with the state-of-the-art models for using real transcriptomic data and WSI patches. For TCGA-GBMLGG, the C Index evaluated for GSCNN Mobadersany et al. [2018b] is 0.781, DeepAttnMISL Yao et al. [2020] is 0.734, MCAT Chen et al. [2021] is 0.817, Pathomic fusion Chen et al. [2020] is 0.826, whereas our methodology yields a C Index of 0.866 using real transcriptomic data and a C Index of 0.861 using synthesized transcriptomic data for survival risk estimation. PathGen-X Krishna et al. [2024] obtains a C-Index of 0.81 on the TCGA-GBM cohort; however, their inference stage is unimodal and only considers a section of patients (Grade IV). For gradation, Pathomic fusion Chen et al. [2020] has an AUC of 0.906, whereas our methodology has an AUC of 0.907 on real transcriptomic data and an AUC of 0.890 on synthesized transcriptomic data. For TCGA-KIRC, the C Index evaluated for Pathomic fusion Chen et al. [2020] is 0.720 whereas for our methodology is 0.697 on real transcriptomic data and 0.681 on synthesized transcriptomic data. However, Pathomic fusion uses ROIs annotated by expert pathologists for predicting grades, while we use transcriptomic features to co-attend to the WSI patch embeddings without requiring any expert intervention, thus reducing expert annotation overheads. For TCGA-UCEC, DeepSetsZaheer et al. [2017], DeepAttnMISLYao et al. [2020], and MCATChen et al. [2021] achieve a C Index of 0.593, 0.586, and 0.622, respectively, whereas our methodology yields a C Index of 0.680 and 0.673 using real and synthesized transcriptomes, respectively, for survival risk estimation. On TCGA-BRCA, our methodology achieved a C-index of 0.720 for both rel and synthesized transcriptomic profiles, compared to 0.522, 0.577, and 0.580 for DeepSetsZaheer et al. [2017], DeepAttnMISLYao et al. [2020], and MCATChen et al. [2021], respectively. We note that the considered SOTA models, DeepSets, DeepAttnMISL, and MCAT, do not perform gradation, so a direct comparison on grade prediction was not possible. For similar reasons, we could only compare the performance of our model with Pathomic fusion for TCGA-GBMLGG gradation and TCGA-KIRC survival risk estimation. Since our model performs better or at least comparably with the state-of-the-art models using real transcriptomic data, and we observe that the predictive performance between using real or synthesized transcriptomic data is very close with no significant difference for the test set population (Wilcoxon rank-sum test $p$-value >0.05>0.05, Supplementary Table LABEL:s-result_main), we may conclude that we achieve state-of-the-art performance using gene expression levels synthesized by our model, PathGen.
Transparency evaluation for clinical interpretability
@@
To visualize intra-tumour heterogeneity, we use our trained model to predict grade and survival risk for individual WSI patches independently. The distributed prediction allows the identification of regions of varied grades within a WSI, highlighting the heterogeneity for clinical assistance (Figures 3.b, 4.b and Supplementary Figures LABEL:s-fig_gbmlgg1.b, LABEL:s-fig_kirc1.b, LABEL:s-fig_cptac_ucec.b). It is evident that higher grade regions are correlated to higher risks, as is expected (Figure 4.e). By observing distributed predictions at different magnification levels corresponding to a WSI, we find that using WSIs at higher resolution helps to obtain more precise predictions for tissue regions, as is expected (Supplementary Figure LABEL:s-fig_mag).
@@
Further, we consider the aggregate of the heterogeneous (distributed) predictions and compare it with the grade predicted by the model, considering all WSI patches as input (non-distributed prediction). For the TCGA cohorts, the distributed gradation performance is not significantly different from a one-time prediction with all WSI patches (Wilcoxon ranksum test $p$-value >> 0.05, Figure 5.a.i). Thus, we see from the results that there is indeed a distributed variation of grade when predicted locally for patches and mapped onto the WSIs as a heatmap, both within and across cohorts, which may provide an additional visual presentation of tumour heterogeneity to the clinician expert (pathologist) to choose whether or not to draw any further intuition from them.